### v3

# The economics of global personality diversity

Abstract
This study explores the relationship between personality diversity and national economic performance, introducing the Global Personality Diversity Index ($\Psi$-GPDI) as a novel metric. Leveraging a dataset of 760,242 individuals across 135 countries, we quantify within-country diversity based on the Big Five personality traits. Our findings reveal that personality diversity accounts for 19.9% of the variance in GDP per capita and provides an additional 2.8% explanatory power beyond institutional quality and immigration, underscoring its unique contribution to economic vitality. Through multi-factor analysis, we demonstrate how personality diversity complements existing economic frameworks, offering actionable insights for policymakers seeking to enhance innovation, productivity, and resilience. This research positions psychological diversity as a critical yet under explored factor in driving economic growth, bridging the fields of psychology and economics.
Introduction
Team diversity enhances performance by fostering innovation and improving problem-solving across various fields. Studies indicate that heterogeneous groups outperform homogeneous ones due to their diverse perspectives. For instance, a comprehensive analysis of over 9 million collaboratively authored research papers involving about 6 million scientists found that diversity in ethnicity, age, gender, and institutional affiliation all positively correlate with research impact, with ethnic diversity leading to up to 40% more citations [1].
While the term “diversity” today often emphasises categorical group differences such as ethnicity, gender, and culture, this approach can sometimes overlook the critical variations within these categories, potentially implying homogeneity that does not exist. Individual differences, encompassing both personality and intellectual abilities, represent a more fundamental level of diversity. By focusing on personality diversity, this study explores a dimension of human variability that transcends group categorizations, reaching the essence of what it means to be diverse.
The significance of personality diversity is evident not only in research teams but also in high-growth firms. A large-scale study involving over 21,000 startups found that those with diverse founder personalities were 8–12 times more likely to succeed, underscoring the role of psychological differences in innovation and economic performance [2].
Similarly, diversity within corporate boards has been shown to significantly enhance innovation at the organizational level. An analysis of 4,448 firms and 43,639 directors found that companies with more diverse boards not only generated a higher volume of patents but also developed innovations of greater impact, as indicated by increased citation frequencies. This highlights the transformative role diverse expertise and perspectives play in fostering groundbreaking ideas and advancing organisational success[3].
While the advantages of diversity are well-established at the micro level—among research teams, startups, and corporations—the broader impact of diversity on economic performance at the macro level remains insufficiently explored. Understanding income disparities between countries is a fundamental aspect of economic research. Acemoglu, Johnson, and Robinson[4], whose work earned them the 2024 Nobel Prize, highlighted how inclusive institutions promote prosperity, whereas extractive institutions hinder economic growth. Nonetheless, institutional factors alone do not fully explain these income disparities, underscoring the need to examine additional dimensions that influence economic outcomes.
Beyond institutional frameworks, immigration emerges as a second significant factor influencing economic performance through various forms of diversity, including birthplace diversity[5], ethnic diversity[6], and genetic variation[7]. The "Out of Africa" theory suggests a non-monotonic relationship between genetic diversity and economic performance, where excessive diversity contributes to Africa’s economic challenges[8], insufficient diversity is linked to South America’s persistent issues[7], and regions like Asia and Europe maintain an optimal "Goldilocks" level of diversity that fosters economic stability and growth.
A third, underexplored dimension is personality. Recent research by Shleifer and colleagues[9, 10, 11, 12, 13, 14, 15, 16] aims to delve into individual cognitive processes and their implications for decision-making. Similarly, Akerlof, Holden, and Li[17] have developed models of human cognition and reasoning. And on the economic geography front, Garretsen et al have shown links between regional economic growth and the presence of specific personality features such Conscientiousness[18]. However, the relationship between individual-level personality traits and aggregate macroeconomic outcomes has received limited attention in economic studies to date.
Building on the exploration of individual personality traits, a crucial yet as yet unexamined dimension is national personality diversity. While studies like McCrae and Terracciano [19] have focused on differences in average personality traits between countries to understand cross-cultural variations, they did not examine the economic implications of these traits nor the variability of personality within nations. This gap makes the case for exploring how within-nation personality diversity, rather than just average national traits, might influence economic performance.
We hypothesize that personality diversity could enhance a nation’s innovation and economic growth. Countries with diverse personality traits may better handle challenges and foster creativity. This concept is reflected in the impact of immigrants who drive economic development. For example, Ugur Sahin and Özlem Türeci, Turkish immigrants to Germany, founded BioNTech, developing a leading COVID-19 vaccine. Elon Musk, who migrated from South Africa to the U.S., revolutionised industries with Tesla and SpaceX. Sergey Brin, an immigrant from the Soviet Union, co-founded Google. These entrepreneurs introduced transformative leadership, profoundly impacting economies.
Building on these observations, our study investigates whether national personality diversity correlates with higher GDP per person employed. We posit that psychologically diverse populations contribute a wider array of perspectives, skills and problem-solving approaches, thereby enhancing economic performance.
To test this hypothesis, we conducted a multifactor analysis incorporating known variables influencing national income: immigration rates and expropriation risk (institutional strength proxy), as highlighted by Acemoglu et al. [4] Including these factors, we aim to show that personality diversity explains additional variance in economic performance. Inspired by the Fama-French three-factor model, we seek to identify factors that complement known measures, particularly institutional strength and immigration.
We introduce the Global Personality Diversity Index ($\Psi$-GPDI), a metric quantifying national personality diversity, constructed from data on 760,242 individuals across 135 countries using the Big Five personality traits [20]. Calculating trait variability within each country provides a standardised measure, allowing examination of correlations between $\Psi$-GPDI and economic indicators.
Understanding the impact of national personality diversity has significant implications for economic policy. If personality diversity contributes to economic performance, policymakers could foster psychological diversity through education, workforce development, and social initiatives—moving the conversation beyond just attracting skilled immigrants. This perspective adds a new dimension to the factors driving economic growth, complementing institutional and cultural theories. By bridging psychology and economics, our study provides a nuanced understanding of how individual differences influence macroeconomic outcomes, shifting the focus from average personality profiles to within-population diversity.
In the following sections, we present our findings on the correlations between personality diversity and economic indicators, followed by a discussion of their implications, limitations, and significance. The detailed methodologies used for constructing the $\Psi$-GPDI and sourcing economic data are provided in the Methods section, with additional information and figures available in the supplementary materials.
Results
To explore global personality diversity, we analysed a comprehensive dataset comprising responses from 760,242 individuals across 135 countries collected between 2005 and 2023. Each participant completed the 300-item IPIP-NEO personality inventory, providing detailed assessments of the Big Five personality traits and their 30 facets. The dataset includes basic demographic information such as country of residence, age, and gender, allowing us to examine personality patterns across different nations and demographic groups.
Global patterns in personality diversity
Personality diversity varies significantly across countries, as revealed through hierarchical clustering of detailed 30-dimensional personality data from respondents in each country (Figure 1). This analysis, using machine learning, uncovers clear and striking differences in personality trait combinations across nations, offering a compelling visualisation of global diversity.
Expected clusters emerge among countries with strong cultural, historical, or geographical ties. For instance, New World British colonies such as Canada, Australia, and Ireland cluster closely together, reflecting shared heritage and similar societal influences. Neighbouring European countries, including Spain and Portugal, as well as Austria and Germany, also align, showcasing the role of geographic proximity and shared cultural backgrounds. Similarly, Nordic countries—Denmark, Norway, and Sweden—form a distinct cluster, consistent with their intertwined histories and cultural cohesion.
Regional patterns are also evident across other parts of the world. In Latin America, nations such as Argentina and Uruguay, along with Colombia and Venezuela, form cohesive clusters, reflecting strong regional ties and shared cultural heritage. East Asian countries, including South Korea and Japan, cluster closely, indicative of shared cultural influences. Similarly, Caribbean nations like Jamaica, Trinidad and Tobago, and Barbados group together, highlighting regional similarities in personality profiles.
In addition to these familiar groupings, the clustering uncovers surprising pairings that suggest intriguing underlying connections. Serbia and Egypt, for example, as well as Israel and Georgia, cluster together, revealing similarities in personality trait combinations that transcend their geographical and historical contexts. These unexpected alignments may point to deeper sociocultural or historical dynamics worth exploring further.
Wealthier nations, such as the United States and many Western European countries, often cluster together, hinting at possible links between personality traits and economic development. Countries with significant immigration histories, including the United States and Canada, show patterns reflecting their diverse and multifaceted personality compositions. These findings demonstrate how hierarchical clustering reveals both expected and surprising patterns, shedding light on how geography, culture, history, and socio-economic factors shape national personality profiles.
Figure 1: |||| Hierarchical clustering of Big Five personality facets across 135 countries. The dendrogram represents the clustering of countries based on median scores across 30 Big Five personality dimensions, using data from a global sample (n = 760,242). The analysis reveals significant variation in personality profiles across countries, with patterns linked to shared cultural, economic, historical, and geographic contexts. While familiar regional pairings are evident, the clustering also highlights unexpected and intriguing combinations, suggesting complex interplays between personality and national characteristics.
Personality diversity across countries
Personality diversity varies significantly across countries, as revealed through an analysis of pairwise personality distances among individuals in each nation (Figure 3). These distances, calculated using cosine similarity on 30-dimensional personality vectors derived from the IPIP-NEO inventory, provide an exhaustive measure of diversity. With hundreds of thousands of individuals in the dataset, the combinations of pairwise comparisons are immense, offering a uniquely high-resolution view of actual personality diversity within each country. This approach allows us to uncover nuanced patterns that go beyond traditional demographic or cultural assessments.
Some nations exhibit single, sharp peaks in their distributions, reflecting relatively uniform personality traits across their populations. In contrast, others display multimodal distributions, suggesting distinct personality clusters. For instance, New Zealand shows a multimodal pattern, likely influenced by differences between Māori and non-Māori populations. South Africa also displays multiple peaks, reflecting its diverse ethnic and linguistic groups, while India’s diversity likely arises from its vast regional, cultural, and religious variety. Belgium’s pattern may reflect the linguistic and cultural divide between Flemish- and French-speaking communities.
In North America, the United States and Canada exhibit broader, multimodal patterns, likely shaped by their immigrant histories and the presence of indigenous populations. Conversely, smaller or more culturally homogenous countries often display single peaks, indicating greater uniformity in personality traits. The variation in medians across countries further highlights differing levels of personality diversity, with nations characterised by high immigration or cultural heterogeneity often demonstrating greater diversity.
These results provide an unprecedentedly detailed view of personality diversity, offering insights into the cultural, historical, and socio-demographic factors that shape national profiles. The high-resolution analysis opens new avenues for understanding the role of personality diversity in social and economic outcomes globally.
In-country personality diversity
In-country personality diversity is evident in the hierarchical clustering analysis of the United States, shown as an example in Figure 2. The clustering reveals natural groupings of individuals with similar patterns of traits across 30 personality dimensions. This analysis was conducted on a randomly sampled subset representing 10% of the total U.S. dataset (n=446,783).
Figure 2: |||| In-country personality combinations cluster into distinct functional types with economic significance. This dendrogram and heatmap illustrate eight personality clusters (Pathfinder, Leader, Advisor, Translator, Yearner, Pioneer, Strategist, Uniter) based on Big Five personality traits across 30 dimensions (n = 44,678), using the United States as an example. The clusters reveal the structure of the PLATYPUS model, showing clear differences in personality profiles associated with varying economic roles and capabilities. Entrepreneurial personalities are notably concentrated in the Pathfinder and Leader clusters, which are linked to founders of startups, innovative and and high-growth firms. Supplementary Information provides the distribution of predicted entrepreneurial personalities by cluster.
The hierarchical clustering identified eight optimal clusters, though preliminary analysis also considered a 13-cluster solution (see Methods Methods for details). This is consistent with recent research suggesting a typology of 13 distinct personality trait combinations [21]. However, the eight-cluster solution was selected due to stronger natural clustering signals in the data.
The PLATYPUS Model of Economic Personality Diversity
Type Archetype Description Key Big Five Personality Facets Economic Functions Pathfinder Proactive Collaborators High in Agreeableness (Trust); Openness (Adventurousness) and Extraversion (Activity Level) Build cohesive teams and foster an inclusive, collaborative culture. Leader Charismatic Leaders High in Agreeableness (Co-ooperation and Morality) and Conscientiousness (Dutifulness). Foster collaboration and drive sustainable growth. Advisor Empathetic Guides High in Conscientiousness (Dutifulness) and Agreeableness (Morality and Altruism) Maintain social infrastructure, essential for productivity. Translator Optimistic Connectors High in Extraversion (Cheerfulness, Excitement-Seeking, Gregariousness) Creating cohesive social and professional networks. Yearner Emotional Idealists High in Openness (Imagination, Emotionality) and Emotional Stability (Anxiety) Offering unique insights into human experiences. Yearner Creative Empaths High in Openness (Artistic Interests) and Low Conscientousness (Cautiousness) Foster innovation and social well-being through creativitiy and spontaneity. Uniter Balanced Dependables Moderate in all Big Five traits, showing versatility and flexibility Bridge gaps between creativity and practical implementation. Strategist Practical Realists High in Conscientiousness (Orderliness, Cautiousness), Low in Agreeableness (Sympathy) Thriving in roles that demand precision and careful planning.
Table 1: Eight distinct personality types were identified within the United States (n = 44,678) using Big Five personality traits. Each type is described by an archetype, key personality facets, and its economic function, illustrating the diverse roles individuals play in fostering innovation, collaboration, and productivity.
To examine the functional implications of these clusters, we applied a global Big Five 30-facet machine learning-based personality predictor with 88% accuracy, developed in a prior study (The Impact of Founder Personality on Startup Success). This model was used to predict the percentage of individuals within each cluster likely to possess personality traits associated with successful entrepreneurs. The results reveal that two of the eight clusters have significantly higher proportions (approximately 15%) of individuals predicted to exhibit entrepreneurial traits. These findings suggest that the typology of personality clusters not only captures distinct psychological profiles but also has meaningful economic implications, with certain clusters displaying a higher propensity for entrepreneurship.
Figure 3: |||| Global Personality Diversity Index ($\Psi$-GPDI) distributions across 135 countries. Ridgeline plot illustrating the Global Personality Diversity Index ($\Psi$-GPDI) for 135 countries, based on pairwise cosine similarity distances in personality traits. The $\Psi$-GPDI reveals the extent of personality diversity within each country, showcasing variations in median values, distribution shapes, and multimodal patterns. These findings highlight distinct national personality dynamics, with implications for cultural cohesion, innovation potential, and economic outcomes.
The Global Personality Diversity Index
To capture and quantify in-country personality diversity, we developed the Global Personality Diversity Index ($\Psi$-GPDI), a measure based on the pairwise cosine similarity of 30-dimensional personality trait vectors derived from the IPIP-NEO inventory. The GPDI reflects the extent of variation in personality within each country, providing a high-resolution view of how personality traits are distributed across populations. By incorporating pairwise comparisons of hundreds of thousands of individuals, the GPDI offers a robust and exhaustive assessment of personality diversity at the national level.
Cross-Country Variations in Personality Diversity The distributions of the $\Psi$-GPDI vary significantly between countries, as visualised in Figure 1b. Some nations, such as New Zealand and South Africa, exhibit multimodal distributions, indicating the presence of distinct personality subgroups within their populations. These patterns likely reflect the influence of cultural, historical, or ethnic diversity, such as the distinct contributions of Māori and non-Māori populations in New Zealand or the varied ethnic and linguistic groups in South Africa. In contrast, countries like Japan and Finland show single, sharp peaks, suggesting greater homogeneity in personality traits.
The medians and ranges of the $\Psi$-GPDI also vary across nations, with countries characterised by high immigration, such as the United States and Canada, showing broader and more diverse distributions. These findings highlight how factors like cultural mixing, historical legacies, and societal structures influence in-country personality diversity.
Statistical Comparisons of Distributions Pairwise Kolmogorov-Smirnov (KS) tests, detailed in Supplementary Figure S2, reveal that the vast majority of countries differ significantly in their GPDI distributions. These results underscore the uniqueness of each nation’s personality diversity profile, emphasising the value of examining personality at the country level. The significant differences observed in these tests reinforce the robustness of the GPDI as a tool for comparing personality diversity across global populations.
Together, the GPDI and the cross-country analyses provide an unprecedented view of global personality diversity, uncovering patterns that offer insights into cultural, economic, and historical factors shaping personality variation. This framework opens new avenues for understanding the implications of personality diversity for social and economic outcomes.
GPDI correlates with national income The analysis reveals a significant relationship between the Global Personality Diversity Index ($\Psi$-GPDI) and national income across 135 countries, measured as GDP per person employed (Figure 4). At moderate levels of GPDI (0.27–0.31), a strong positive correlation is observed, with high-income nations such as Luxembourg, Ireland, and the United States achieving particularly strong economic performance. These results reveal that, in addition to well-known drivers of GDP variance such as immigration and institutional structures, personality diversity itself is a measurable and significant factor influencing national economic outcomes across a broad global sample.
However, as the GPDI exceeds 0.31, the relationship plateaus or slightly declines, indicating there may be diminishing returns to diversity. This pattern suggests that beyond a certain threshold, the benefits of diversity may be tempered by other societal or structural factors.
Figure 4: |||| GDP per capita and Personality Diversity Index ($\Psi$-GPDI) across 135 countries. Scatter plot showing the relationship between the Personality Diversity Index ($\Psi$-GPDI) and GDP per capita (in constant 2005–2023 PPP \$) for 135 countries. A positive association is observed, with higher personality diversity correlating with greater economic productivity in most regions. However, deviations such as a plateau or decline in some cases highlight the complex interplay between diversity and economic outcomes. Labels identify select OECD and G20 countries. This analysis underscores the potential role of personality diversity in driving national income and innovation metrics.
Statistical analysis reinforces these findings, with linear and parabolic models explaining 13.9% of the variance in GDP per person and 10.3% in GNI per capita (R² values; ***p < 0.001). The parabolic fit captures the diminishing returns seen in countries like Angola, where high personality diversity does not correspond to higher economic performance.
Multifactor model results An initial experimental multifactor model was developed to predict GDP per person employed (constant 2021 PPP), incorporating a range of country-level economic features alongside Personality Diversity. These features included exports and imports of goods and services, research and development expenditure, the number of researchers per million people, and measures of migration such as the share of migrants and skilled migrants.
In the Null Model, which excluded Personality Diversity, the adjusted R-squared value was 0.7358, indicating that approximately 73.58% of the variability in GDP per person employed could be explained by the selected economic factors.
When Personality Diversity was added in the Personality Model, the adjusted R-squared value increased to 0.7638, explaining 76.38% of the variability in GDP per person employed. This represents an improvement of around 2.8% in the model’s explanatory power. Furthermore, Personality Diversity was found to have a significant positive coefficient (15.3458, p = 0.00148), demonstrating that it is a statistically significant predictor of GDP per person employed.
These results suggest that Personality Diversity provides a meaningful contribution to explaining national economic outcomes, even when controlling for other key economic factors. This finding highlights the potential importance of psychological and cultural variables in economic modelling, offering new perspectives on the drivers of economic productivity.
Figure 5: |||| Multifactor analysis identifying Personality Diversity Index ($\Psi$-GPDI) as a key third factor in explaining national income. Scatterplots and correlation matrix illustrating the relationships between GDP per capita, Personality Diversity Index ($\Psi$-GPDI), migration, and expropriation risk. The analysis positions $\Psi$-GPDI as a third key factor, alongside two established factors—migration (Corr = 0.488) and expropriation risk (Corr = 0.716)—in explaining variance in national income. The moderate correlation between $\Psi$-GPDI and GDP (Corr = 0.453) highlights its potential as a novel explanatory variable in understanding economic outcomes, complementing demographic and economic risk factors.
Table 2: |||| Contributions of Expropriation, Migration, and Personality to GDP, showing the adjusted R-squared values for individual and combined factors. The table also includes Variance Inflation Factor (VIF) values for each factor, assessing multicollinearity. The associated Venn diagram illustrates the overlapping contributions of these factors to national income, highlighting the links between personality diversity, migration, expropriation risk, and GDP.
Models Adjusted R-squared
GDP   Expropriation risk 50.8% (***)
GDP   Immigration 23.1% (***)
GDP   $\Psi$-GPDI 19.9% (***)
GDP   Expropriation risk + Immigration 57.1% (***)
GDP   Expropriation risk + $\Psi$-GPDI 56.1% (***)
GDP   Immigration + $\Psi$-GPDI 40.0% (***)
GDP   Expropriation risk + Immigration + $\Psi$-GPDI 62.8% (***)
Discussion
Geographic variation in personality has been extensively studied, revealing significant differences both within and between countries. For instance, studies of up to 56 different countries [22, 19, 23] have documented considerable variation in Big Five personality traits. Within countries, regional personality differences have been observed, including three distinct personality zones in the United States by region [24], variations shaped by topography and ecology in the United States [25], and geographic clustering of traits like neuroticism and conscientiousness in the United Kingdom [18].
Previous efforts to cluster personality by country [23], based on Big Five domains across 36 nations, identified patterns that partially align with our findings. However, by leveraging 30 dimensions of personality data for each country, our clustering approach provides a higher-resolution analysis of personality similarities between nations. Using hierarchical clustering on data from 135 countries, we observe notable personality variation across nations, visualised in a spherical hierarchical dendrogram (Figure 1). This approach offers deeper insights into global personality diversity.
Global clusters of personality
At the highest level, three primary clusters emerge, reflecting broad personality groupings:
  • Green Cluster: This group predominantly comprises advanced Western economies, including the United States, United Kingdom, Australia, Canada, and New Zealand. These countries share cultural, economic, and historical ties, often characterised by similar personality traits.
  • Blue Cluster: Representing many Eastern European, Central Asian, and smaller nations, this group includes Ukraine, Romania, Bulgaria, Kazakhstan, Cyprus, and Georgia. These nations exhibit shared traits that may reflect historical and socio-political influences within the region.
  • Red Cluster: Encompassing countries from the Global South, such as Kenya, Tanzania, Botswana, Brazil, Colombia, and Chile, this cluster highlights shared colonial histories and developmental trajectories that appear to shape collective personality profiles.
Patterns and anomalies in personality pairings
At the finer levels of clustering, expected patterns emerge, with geographic neighbours sharing similar historical and cultural roots often appearing adjacent on the dendrogram. For instance, Brazil and Chile, Spain and Portugal, and Austria and Germany form intuitive pairings.
However, some surprising pairings also surface, such as New Zealand with Estonia, Paraguay with Croatia, and Greece with Argentina. These unexpected alignments may reveal latent similarities in worldview or shared behavioural tendencies, offering a new lens through which to understand the underlying drivers of personality variation across nations. Such pairings invite further exploration into how geography, history, and shared experiences influence collective personality traits.
The significance of national personality profiles
The personality traits of a country’s population offer insights far beyond cultural stereotypes, as they often deviate significantly from national caricatures [26]. Importantly, personality traits are tied to economic outcomes, with research demonstrating their predictive power for national and regional prosperity. For instance, traits such as extraversion and openness, when combined with intelligence, are strongly associated with GDP per capita and economic success [27]. Similarly, regional patterns of traits like neuroticism and conscientiousness influence urban economic growth, while broader cultural and psychological profiles shape productivity differences across regions [18, 24].
On a global scale, economic preferences like patience have been shown to explain 80% of the variance in GDP [28]. While not a personality trait itself, patience has been linked to traits such as conscientiousness and self-discipline in large-scale studies, further highlighting the connection between personality and economic outcomes [29].
Personality diversity and economic success
While individual personality traits play a critical role in predicting success, it is the diversity of personality types working in concert that drives the most significant outcomes. Research on 21,187 startups, for example, revealed that while specific traits are linked to entrepreneurial success, teams with a diverse mix of personality types are 8–12 times more likely to succeed [2]. This underscores the importance of a broad range of traits across individuals for achieving collective goals.
At the national level, this principle extends naturally. A flourishing economy requires a variety of roles suited to different personality profiles. Entrepreneurs and innovators are vital, but so are lawmakers, regulators, educators, scientists, and other professionals. Each of these roles is likely to align with distinct psychological traits. Indeed, a large-scale study involving 128,279 individuals across 3,513 occupations confirmed that distinct personality profiles strongly align with specific job roles [30]. This contemporary study aligns with a long literature of personality linkages to occupational roles that dates back over a century with EK Strong’s ideas[31] conceptualised in John L. Holland’s personality-occupational types and his classic book: Making Vocational Choices[32] and later work with Costa and McCrae[33] that explored connections between Big Five and his hexagonal model. While individuals in similar roles often exhibit similar traits, those in different roles display diverse personality footprints, reflecting the range of talents required for societal and economic success.
Personality Clusters Within Nations
Our data reveal the rich diversity of personality clusters within countries, using the United States as an example. We identified eight distinct clusters, each representing groups of individuals with similar psychological traits (Figure 2).
These clusters are labelled as PLATYPUS to reflect their unique characteristics: Pathfinder (Proactive Collaborators), Leader (Charismatic Leaders), Advisor (Empathetic Guides), Translator (Optimistic Connectors), Yearner (Emotional Idealists), Pioneer (Creative Empaths), Uniter (Balanced Dependables), and Strategist (Practical Realists) (Table 1).
Among these, two clusters—Pathfinders and Leaders—stood out for their high proportion of individuals predicted to be entrepreneurs, based on a model trained on data from thousands of entrepreneurs in previous research [2]. While individuals in these clusters may be particularly suited to entrepreneurial roles, others align more closely with equally essential roles, such as health professionals, scientists, educators, and other critical contributors to a thriving and balanced economy. A flourishing economy depends not only on entrepreneurship but also on a diversity of roles that collectively drive innovation, resilience, and growth.
Global Personality Diversity Index scores
We calculated the Global Personality Diversity Index ($\Psi$-GPDI) scores for each country using high-performance computing. These calculations involved pairwise comparisons of personality footprint distances across 30 dimensions for 760k individuals from 135 countries. This process was computationally intensive, involving over 1.5 billion combinations, ensuring an analysis that is both exhaustive and highly fine-grained.
The results are illustrated in Figure 3, which highlights the distributions of personality diversity within each country. The figure showcases key variations, including differences in median values, the shapes of the distributions, and multimodal patterns. These insights provide a detailed view of how personality diversity manifests globally, offering a nuanced understanding of the unique psychological profiles within nations.
Among the most multimodal countries in the joy plot are Turkey, New Zealand, and Brazil, reflecting distinct subgroups within their populations. Turkey’s diversity likely arises from its cultural position bridging Europe and Asia, as well as urban-rural contrasts. New Zealand showcases differences between indigenous Māori communities and the broader population, while Brazil reflects significant socio-economic stratification and regional cultural variation. These countries exemplify how historical, cultural, and demographic factors contribute to complex personality profiles.
We also compared the similarity of each country’s GDPI distribution and found nearly all countries have distinct personality diversity distributions, as shown by pairwise Kolmogorov-Smirnov tests (Supplementary Figure 5).
GPDI Validates global personality patterns
Our GPDI is also self-report-based, aligning methodologically with Schmitt’s dataset, often considered the gold standard in personality assessment. To further validate GPDI, we compared its scores against McCrae et al. [19], which utilised observer ratings from 12,156 participants across 51 countries. While observer ratings differ from self-reports, they offer an independent benchmark that minimises biases such as cultural response styles. A statistically significant correlation (p = 0.011) between McCrae’s standard deviations and GPDI scores underscores the robustness of our index and its consistency across methodologies. These findings highlight the validity of GPDI in capturing global patterns of personality diversity, leveraging complementary data sources to provide deeper insights into cross-national differences.
Limitations
It is important to recognize that while a country may exhibit a personality-diverse population, the economic benefits of such diversity might remain unrealized if individuals with similar personality traits primarily associate with each other. Research by Ma et al. [30] highlights how "social sorting," driven by individuals with lower perceived personal control, can constrain the potential advantages of diversity. External factors, such as political unrest, economic instability, or governance structures like totalitarian regimes, may further exacerbate low personal control, leading to increased similarity attraction and homogeneity within groups. This dynamic underscores the importance of fostering cross-group interactions and addressing structural influences to unlock the full economic and social potential of a personality-diverse society.
Another limitation of this study is the potential for sampling bias, as the data were collected voluntarily over more than two decades from 760,242 participants, all responding to an English-language version of the IPIP-NEO inventory. The age distribution of the sample skews younger, with a median age of 22 compared to the global median age of 28.2 in 2010 (UN data). However, the decline in participation with age in the sample broadly mirrors global population trends, suggesting the pattern may partly reflect demographic realities. Despite these limitations, the sample spans a broad age range, capturing responses from individuals aged 10 to 99, and reflects the inventory’s enduring popularity over decades.
Importantly, the distribution of the sample by age and by country aligns well with global population distributions, exhibiting statistically significant correlations with UN 2010 data. The age distribution of the Johnson sample shows a strong linear correlation (r = 0.64, p < 0.0001), as illustrated in Extended Data Fig. 1, supporting its representativeness. Similarly, the country distribution of the sample also exhibits a significant linear correlation (r = 0.54, p < 0.001), as shown in Extended Data Figure 2. These consistent correlations highlight the broad alignment of the sample with global population patterns, reinforcing the robustness of the dataset for cross-national analysis.
Significance
The relationship between the Personality Diversity Index ($\Psi$-GPDI) and GDP per capita reveals a positive association across 135 countries (Figure 4), with higher personality diversity generally correlating with greater economic productivity. This trend is supported by a linear fit (R² = 20.87%) and further clarified by a parabolic fit (R² = 24.41%), which highlights deviations in some regions where the economic benefits of diversity appear to plateau or even decline. Such patterns suggest a complex interplay between diversity and economic outcomes, influenced by factors like social integration and institutional structures. Notably, OECD and G20 countries dominate the upper end of the distribution, underscoring the potential of personality diversity to drive national income and innovation metrics while also highlighting the need to address barriers to realising these benefits.
In the Supplementary Information, we present two additional charts focusing on the relationship between the Global Personality Diversity Index ($\Psi$-GPDI) and GDP per capita for OECD + G20 countries and African countries. Both charts reveal similar trends to the primary analysis, with higher personality diversity correlating positively with greater economic productivity. For the OECD + G20 group, the linear and parabolic models demonstrate stronger fits (R² = 32.47% and 41.53%, respectively), highlighting a nuanced relationship where gains from diversity may plateau at higher levels. Similarly, the analysis of African countries (n=24) shows significant positive correlations, underscoring that the link between diversity and economic outcomes persists across vastly different economic and regional contexts. These supplementary analyses provide robust support for the role of personality diversity in influencing national economic performance.
The multifactor analysis positions the Personality Diversity Index ($\Psi$-GPDI) as a significant contributor to national income alongside established factors like expropriation risk and migration. Personality diversity shows a moderate positive correlation with GDP (Corr = 0.453), complementing the stronger correlation observed for expropriation risk (Corr = 0.716). When combined, these factors explain up to 62.8% of the variance in GDP, as reflected in the highest adjusted R-squared value. Notably, $\Psi$-GPDI enhances explanatory power when added to models including expropriation risk or migration, highlighting its novel role in economic analysis. These results underscore the importance of incorporating personality diversity into broader frameworks for understanding national economic outcomes.
These results underscore the importance of personality diversity in shaping economic policy and innovation, suggesting a broader perspective on immigration policy. Moving beyond solely skills-based migration to consider personality diversity could enhance human capital by bringing a wider range of traits that foster creativity, collaboration, and resilience. Integrating diverse personality types can stimulate economic growth through varied approaches to problem-solving and innovation, while also strengthening societal adaptability to global challenges. Policies that promote inclusivity, cross-cultural collaboration, and the active integration of diverse individuals into the workforce can unlock the full potential of personality diversity, driving sustained innovation and long-term prosperity.
Acknowledgements
We extend a special thank you to Proessor Daniel Falster, whose patient and kind introduction to the concepts of diversity in evolutionary ecology some years ago sowed the seeds for this study and was key to making this work possible.
We are grateful to Professor Jean Twenge for her shared research and insights on generational personality, which provided valuable context and inspiration for this work. We also acknowledge Professor Eugene Fama, whose foundational work on the three-factor model inspired aspects of our analytical approach, and Professor Daron Acemoglu, whose groundbreaking research on institutions and economic development influenced the broader framing of this study. Their work has been instrumental in shaping the theoretical foundations of this research.
Author contributions
All authors designed research, analysed data, and undertook investigation. P.X.M. conceptualised the study and led methodology design, project administration, and manuscript writing. X.G. processed data, performed statistical modelling, and contributed to the machine learning framework, while collecting and tabulating data, creating figures, and finalizing artwork. M.C. curated data, created visualizations, and drafted sections of the manuscript. M.A.R. led computational modelling and validated analytical methods. M.L.K. led personality insights, and edited the manuscript. J.A.J. contributed the underlying data and its interpretation, validated personality models, and reviewed drafts. R.H. advised on economic modelling and integrated economic theories. F.B. coordinated interdisciplinary collaboration, integrated contributions, and provided senior authorship oversight. All authors contributed to writing the paper.
Competing interests
The authors have declared that no competing interests exist.
Appendix A Supplementary Information
A.1 Data representativeness
Extended Data Fig. 1: |||| Comparison of Johnson Data Age Distribution and Global Population (UN 2010). The age distribution of the Johnson sample (blue bars) is compared to the global population distribution (red line) based on UN 2010 data. The Johnson sample exhibits a linear correlation (r = 0.64, p = 0.000000001%), supporting its representativeness. A logarithmic transformation is applied to the global population data to align scales and illustrate parallel trends, highlighting the consistency in relative age distribution patterns between the two datasets.
Extended Data Fig. 2: |||| Comparison of Johnson Data Country Distribution and Global Population by Country (UN 2010). The country distribution of the Johnson sample (light blue bars) is compared to the global population distribution by country (red line) based on UN 2010 data, with China and India omitted for clarity. The two datasets show a linear correlation (r = 0.54, p = 0.000007811%), indicating the representativeness of the Johnson sample. The logarithmic trend highlights the differences in scale while preserving the general alignment of patterns across countries.
A.2 Equality test of pairwise cosine similarity distributions of 135 countries
Extended Data Fig. 3: |||| Global personality diversity distribution comparison. The Kolmogorov-Smirnov (KS) test reveals that most countries have significantly different distributions of personality diversity.
A.3 Relationships between PDI and GDP
A                                                                     B
Extended Data Fig. 4: |||| A, OECD + G20 GDP per capita for “Magic 46”. B, GDP per capita only African Countries.
A.4 Entrepreneurship Probability Across US Personality Clusters
Extended Data Fig. 5: The bar chart presents the likelihood of individuals within each US personality cluster (Pathfinder, Leader, Advisor, Translator, Yearner, Pioneer, Strategist, Uniter) exhibiting entrepreneurial traits. These probabilities were derived using a machine learning model based on Big Five personality traits and 30 facets, trained on methods described in McCarthy et al (2023) [2]. The Pathfinder and Leader clusters show the highest entrepreneurial probabilities (2x average probability) emphasizing their strong association with innovation and high growth firms. Other clusters are likely aligned with different and yet also important economic functions. Error bars indicate confidence intervals for the estimates, highlighting the functional diversity revealed by the PLATYPUS framework.

### v4

@@
 These findings demonstrate how hierarchical clustering reveals both expected and surprising patterns, shedding light on how geography, culture, history, and socio-economic factors shape national personality profiles.
-Figure 1: |||| Hierarchical clustering of Big Five personality facets across 135 countries.
+Figure 1: $|$ Hierarchical clustering of Big Five personality facets across 135 countries.
 The dendrogram represents the clustering of countries based on median scores across 30 Big Five personality dimensions, using data from a global sample (n = 760,242).
@@
 dataset (n=446,783).
-Figure 2: |||| In-country personality combinations cluster into distinct functional types with economic significance.
+Figure 2: $|$ In-country personality combinations cluster into distinct functional types with economic significance.
 This dendrogram and heatmap illustrate eight personality clusters (Pathfinder, Leader, Advisor, Translator, Yearner, Pioneer, Strategist, Uniter) based on Big Five personality traits across 30 dimensions (n = 44,678), using the United States as an example.
@@
 The PLATYPUS Model of Economic Personality Diversity
-Type Archetype Description Key Big Five Personality Facets Economic Functions Pathfinder Proactive Collaborators High in Agreeableness (Trust); Openness (Adventurousness) and Extraversion (Activity Level) Build cohesive teams and foster an inclusive, collaborative culture.
-Leader Charismatic Leaders High in Agreeableness (Co-ooperation and Morality) and Conscientiousness (Dutifulness).
-Foster collaboration and drive sustainable growth.
+Type Archetype Description Key Big Five Personality Facets Economic Functions
+Pathfinder Proactive Collaborators High in Agreeableness (Trust); Openness (Adventurousness) and Extraversion (Activity Level) Build cohesive teams and foster an inclusive, collaborative culture.
+Leader Charismatic Leaders High in Agreeableness (Co-ooperation and Morality) and Conscientiousness (Dutifulness). Foster collaboration and drive sustainable growth.
 Advisor Empathetic Guides High in Conscientiousness (Dutifulness) and Agreeableness (Morality and Altruism) Maintain social infrastructure, essential for productivity.
@@
 These findings suggest that the typology of personality clusters not only captures distinct psychological profiles but also has meaningful economic implications, with certain clusters displaying a higher propensity for entrepreneurship.
-Figure 3: |||| Global Personality Diversity Index ($\Psi$-GPDI) distributions across 135 countries.
+Figure 3: $|$ Global Personality Diversity Index ($\Psi$-GPDI) distributions across 135 countries.
 Ridgeline plot illustrating the Global Personality Diversity Index ($\Psi$-GPDI) for 135 countries, based on pairwise cosine similarity distances in personality traits.
@@
 This pattern suggests that beyond a certain threshold, the benefits of diversity may be tempered by other societal or structural factors.
-Figure 4: |||| GDP per capita and Personality Diversity Index ($\Psi$-GPDI) across 135 countries.
+Figure 4: $|$ GDP per capita and Personality Diversity Index ($\Psi$-GPDI) across 135 countries.
 Scatter plot showing the relationship between the Personality Diversity Index ($\Psi$-GPDI) and GDP per capita (in constant 2005–2023 PPP \$) for 135 countries.
@@
 This finding highlights the potential importance of psychological and cultural variables in economic modelling, offering new perspectives on the drivers of economic productivity.
-Figure 5: |||| Multifactor analysis identifying Personality Diversity Index ($\Psi$-GPDI) as a key third factor in explaining national income.
+Figure 5: $|$ Multifactor analysis identifying Personality Diversity Index ($\Psi$-GPDI) as a key third factor in explaining national income.
 Scatterplots and correlation matrix illustrating the relationships between GDP per capita, Personality Diversity Index ($\Psi$-GPDI), migration, and expropriation risk.
@@
 The moderate correlation between $\Psi$-GPDI and GDP (Corr = 0.453) highlights its potential as a novel explanatory variable in understanding economic outcomes, complementing demographic and economic risk factors.
-Table 2: |||| Contributions of Expropriation, Migration, and Personality to GDP, showing the adjusted R-squared values for individual and combined factors.
+Table 2: $|$ Contributions of Expropriation, Migration, and Personality to GDP, showing the adjusted R-squared values for individual and combined factors.
 The table also includes Variance Inflation Factor (VIF) values for each factor, assessing multicollinearity.
@@
 Extended Data Fig.
-1: |||| Comparison of Johnson Data Age Distribution and Global Population (UN 2010).
+1: $|$ Comparison of Johnson Data Age Distribution and Global Population (UN 2010).
 The age distribution of the Johnson sample (blue bars) is compared to the global population distribution (red line) based on UN 2010 data.
@@
 Extended Data Fig.
-2: |||| Comparison of Johnson Data Country Distribution and Global Population by Country (UN 2010).
+2: $|$ Comparison of Johnson Data Country Distribution and Global Population by Country (UN 2010).
 The country distribution of the Johnson sample (light blue bars) is compared to the global population distribution by country (red line) based on UN 2010 data, with China and India omitted for clarity.
@@
 Extended Data Fig.
-3: |||| Global personality diversity distribution comparison.
+3: $|$ Global personality diversity distribution comparison.
 The Kolmogorov-Smirnov (KS) test reveals that most countries have significantly different distributions of personality diversity.
@@
 A                            B
-Extended Data Fig. 4: |||| A, OECD + G20 GDP per capita for “Magic 46”. B, GDP per capita only African Countries.
+Extended Data Fig. 4: $|$ A, OECD + G20 GDP per capita for “Magic 46”. B, GDP per capita only African Countries.
 A.4 Entrepreneurship Probability Across US Personality Clusters